# The impact of pressure oscillations on bubble rising in shear-thinning fluids

### v1

@@ 5.  Conclusions
Using numerical simulations, we investigated the rising dynamics of millimiter-sized bubbles under external pressure oscillations within an incompressible, shear-thinning fluid. Motivated by previous experiments [1], we employed the Carreau–Yasuda constitutive equation to model the shear-thinning properties of the fluid and we considered bubbles with an initial radius between 0.5mm and 2mm. We assumed an axisymmetric geometry and we neglected mass and heat transport between the gas and the fluid. We also neglected the viscosity of the gas, thus avoiding to solve for the flow within the bubble. We employed an isothermal compression law to describe the gas pressure within the bubble.
@@
Before studying the effect of pressure oscillations, we investigated the natural rising motion of the bubbles. We found that bubbles smaller than 1mm closely follow Stokes law. Instead, bubbles approaching a 2 mm radius begin to significantly modify the local viscosity, leading to an increased rising velocity compared to that predicted by Stokes law. We then applied an external pressure driving and studied the bubble rising and shape dynamics. To explore a wide range of cases comparable to those in the experiments of Iwata et al [1], we changed the frequency and the amplitude of the periodic pressure driving.
@@
The external pressure oscillations drive periodic bubble volume oscillations. This results in the appearance of large local deformation rates, which reduce the effective viscosity in the bubble’s surroundings. As a consequence, the bubble drag decreases and its rising velocity increases by several orders of magnitude compared to that in the absence of the external driving. The bubble remained almost spherical in all simulations, exhibiting pronounced nonlinear radial and rising dynamics. A Fourier analysis of the bubble kinematics confirms the presence of strong nonlinearities, which grow with increasing forcing frequency and amplitude. The radial dynamics is characterized mainly by the first multiples of the driving frequency. Instead, the rising velocity displays contributions from many higher-order harmonics, with larger contributions from even harmonics. Post-processing analysis of the dimensionless numbers confirms that the simulations were conducted in viscous regimes. However, the significant reduction of the shear viscosity can lead to effective Womersley and Reynolds numbers close to unity. This suggests that, even if inertial effects are absent during natural bubble rising, they can become relevant as external pressure oscillations are applied.
@@
Comparison with experimental data[1] revealed a good qualitative agreement, particularly in terms of the observed enhancement in rising velocity. The results are not far from those reported by Iwata et. al. but the trend as the oscillation amplitude is increased does not match that observed in the experiments. Notable discrepancies were also found in the bubble shape throughout the oscillation cycle. We speculate that these differences stem from the absence of an elastic component in the current model, which prevents capturing the full viscoelastic behavior of the experimental system. The mismatch in velocity is also attributed to the lack of elastic effects that can slow down the bubble and modify its rising path.
@@ Acknowledgments
This work was supported by the Ramon y Cajal fellowship RYC2021-030948-I and by the PID2022-139803NB-I00 research grant funded by the MICIU/AEI /10.13039/501100011033 and by the EU under the NextGenerationEU/PRTR program. Mario Riccio was supported by the FPI PhD fellowship PRE2023-UZ-10 funded by MICIU/AEI /10.13039/501100011033.

### v2

@@ 5. Conclusions
Using numerical simulations, we investigated the rising dynamics of millimeter-sized bubbles under external pressure oscillations within an incompressible, shear-thinning fluid. Motivated by previous experiments [1], we employed the Carreau–Yasuda constitutive equation to model the shear-thinning properties of the fluid and we considered bubbles with an initial radius between $0.5\,\mathrm{mm}$ and $2\,\mathrm{mm}$. We assumed an axisymmetric geometry and we neglected mass and heat transport between the gas and the fluid. We also neglected the viscosity of the gas, thus avoiding solving for the flow within the bubble. We employed an isothermal compression law to describe the gas pressure within the bubble.
@@
Before studying the effect of pressure oscillations, we investigated the natural rising motion of the bubbles. We found that bubbles smaller than $1\,\mathrm{mm}$ closely follow Stokes law. Instead, bubbles approaching a $2\,\mathrm{mm}$ radius begin to significantly modify the local viscosity, leading to an increased rising velocity compared to that predicted by Stokes law. We then applied an external pressure driving and studied the bubble rising and shape dynamics. To explore a wide range of cases comparable to those in the experiments of Iwata et al. [1], we changed the frequency and the amplitude of the periodic pressure driving. The external pressure oscillations drive periodic bubble volume oscillations. This results in the appearance of large local deformation rates, which reduce the effective viscosity in the bubble’s surroundings. As a consequence, the bubble drag decreases and its rising velocity increases by several orders of magnitude compared to that in the absence of the external driving.
@@
The bubble remained almost spherical in all simulations, exhibiting pronounced nonlinear radial and rising dynamics. A Fourier analysis of the bubble kinematics confirms the presence of strong nonlinearities, which grow with increasing forcing frequency and amplitude. The radial dynamics is characterized mainly by the first multiples of the driving frequency. Instead, the rising velocity displays contributions from many higher-order harmonics, with larger contributions from even harmonics. Post-processing analysis of the dimensionless numbers confirms that the simulations were conducted in viscous regimes. However, the significant reduction of the shear viscosity can lead to effective Womersley and Reynolds numbers close to unity. This suggests that, even if inertial effects are absent during natural bubble rising, they can become relevant as external pressure oscillations are applied.
@@
Comparison with experimental data [1] revealed a good qualitative agreement, particularly in terms of the observed enhancement in rising velocity. The results are not far from those reported by Iwata et al., but the trend as the oscillation amplitude is increased does not match that observed in the experiments. Notable discrepancies were also found in the bubble shape throughout the oscillation cycle. We speculate that these differences stem from the absence of an elastic component in the current model, which prevents capturing the full viscoelastic behavior of the experimental system. The mismatch in velocity is also attributed to the lack of elastic effects that can slow down the bubble and modify its rising path.
The main findings may be summarized by the following points:
  • Pressure oscillations induce strong volumetric bubble oscillations, generating large local strain rates in the surrounding fluid.
  • The shear-thinning effect significantly reduces the effective viscosity near the bubble, leading to pronounced drag reduction.
  • The bubble rising velocity increases by several orders of magnitude compared to natural rising, especially at higher forcing amplitudes and frequencies.
  • Bubble dynamics are highly nonlinear, with the rising velocity exhibiting multiple harmonics and strong deviations from sinusoidal behavior.
  • Smaller bubbles experience the largest relative enhancement in rising speed, as their natural motion does not trigger shear thinning.
  • Numerical simulations qualitatively agree with experiments, but discrepancies in bubble shape and rise velocity suggest the need for viscoelastic constitutive models to fully capture the physical behavior of the reference work.
@@ Acknowledgments
This work was supported by the Ramon y Cajal fellowship RYC2021-030948-I and by the PID2022-139803NB-I00 research grant funded by the MICIU/AEI /10.13039/501100011033 and by the EU under the NextGenerationEU/PRTR program. Mario Riccio was supported by the FPI PhD fellowship PRE2023-UZ-10 funded by MICIU/AEI /10.13039/501100011033. The authors thank Dr. Paula Martínez Lera for the help with the Fourier analysis.